Calculate the total surface area of the rectangular prism with length=7,2m,breadth=5m and height=3,32m
Answer by macston(5194)

.
Surface area=2*(length*breadth)+2*(breadth*height)+2*(height*length)
SA=2(7.2m)(5m)+2(5m)(3.32m)+2(3.32m)(7.2m)
SA=2(36m^2)+2(16.6m^2)+2(23.90m^2)
SA=72m^2+33.2m^2+47.80m^2
SA=153.00m^2
ANSWER: Surface area is 153 square meters.